My daughter has many sisters. She has as many sisters as she has brothers. Each of her brothers has twice as many sisters as brothers. How many sons and daughters do I have?
Answer by ankor@dixie-net.com(22740)

My daughter has many sisters. She has as many sisters as she has brothers.
Each of her brothers has twice as many sisters as brothers.
How many sons and daughters do I have?
:
Let d = no. of daughters
Let s = no. of sons
:
"She has as many sisters as she has brothers", therefore:
s = d - 1
:
"Each of her brothers has twice as many sisters as brothers."
d = 2(s-1)
d = 2s - 2
replace s with (d-1)
d = 2(d-1) - 2
d = 2d - 2 - 2
d - 2d = -4
-d = -4
d = 4 daughters
then
4 - 1 = 3 sons